OpenStudy (anonymous):

acetic acid 0.921% (w/w) 100grams. The acetic acid will be measured by weight. how much is used?

OpenStudy (jhannybean):

What does (w/w) mean?

OpenStudy (anonymous):

weight in weight

OpenStudy (jhannybean):

@aaronq come in here. So im guessing it means : \[\sf \frac{0.921 ~g~C_2H_3O_2}{100~g~comp}\]

OpenStudy (anonymous):

i sent the actual question above. with 4 answers..

OpenStudy (aaronq):

acetic acid 0.921% (w/w) 100 grams. means that there is 0.00921*100 g =0.921 g of acetic acid and 100 g - 0.921 g = 99.079 g of solvent

OpenStudy (jhannybean):

oh i see how it works now

OpenStudy (aaronq):

from that picture posted, it's A and C, because of the units

OpenStudy (aaronq):

okay, coooool

OpenStudy (anonymous):

so its a and c? why?

OpenStudy (aaronq):

because of the units, 0.921 g = 921 mg

OpenStudy (anonymous):

ok. thank you so much.. both of you ..

OpenStudy (anonymous):

Are you willing to help me understand other questions?

OpenStudy (aaronq):

no problem guyses, idk if you're asking me, but if so, you should post in the chem section

OpenStudy (anonymous):

im asking either of of you

OpenStudy (anonymous):

if not, would you really consider this chem?

OpenStudy (aaronq):

yeah

OpenStudy (anonymous):

its actually pharmacy compounding math

OpenStudy (aaronq):

pharmacy is based from chem and bio

OpenStudy (anonymous):

are you willing to help me understand a few other questions ?

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

ok thank you i will post the next question

OpenStudy (aaronq):

kk

OpenStudy (anonymous):

if possible explanations please.

OpenStudy (anonymous):

I am trying to get some of these done before work tomorrow

OpenStudy (anonymous):

along with understanding them

OpenStudy (aaronq):

so, v/v means volume by volume, it's pretty straight forward that in 100 mL of 50 % v/v glycerin solution there is 50 mL of glycerin.

OpenStudy (anonymous):

what about the second question

OpenStudy (anonymous):

i should have just read that one. that was common sense but im old..

OpenStudy (anonymous):

is the answer to 17. 6.25ml?

OpenStudy (aaronq):

sorry, i had to step away from the keyboard. yes, that's the answer. you basically just have to watch the units

OpenStudy (anonymous):

its ok . i am trying to do this, which most has nothing to do with what i do at work. and thank you so so much

OpenStudy (anonymous):

here is another one if you can help thank you

OpenStudy (aaronq):

which one?

OpenStudy (anonymous):

i need to do all 3 questions? so what ever you can help with

OpenStudy (anonymous):

well one is d

OpenStudy (anonymous):

so 2 and 3

OpenStudy (aaronq):

b) is 400 mg 0.02*20 g =0.4 g = 400 mg

OpenStudy (anonymous):

wow thank you.. at least i can see it now

OpenStudy (aaronq):

for c) you add the whole masses and make a dilution \(2\%*20g=x\%*(20~g+30~g+30~g)\rightarrow x\%=0.5\%\)

OpenStudy (aaronq):

lol sorry that sounds to weird, i meant "add all the masses" not "the whole.." lol

OpenStudy (anonymous):

add all the grams

OpenStudy (aaronq):

yep

OpenStudy (aaronq):

grams is a unit of mass

OpenStudy (anonymous):

you are a great help.. can you help with a few more? i dont want to over welcome my help

OpenStudy (aaronq):

I would but i have to go do laundry and stuff. Hopefully someone else can come help ya, if you can't find help here, try the chem section.

OpenStudy (anonymous):

ok thank you so much